# CookAR: Affordance Augmentations in Wearable AR to Support Kitchen Tool Interactions for People with Low Vision

### v1

@@ 5.4. Analysis
We recorded participants’ quotes using Zoom. Transcriptions were first done by the video conferencing software, then the research team manually revised the transcripts. We collected 346 distinct quotes across our 10 LV participants, which we analyzed using reflexive thematic analysis (Braun and Clarke, 2006, 2019). The first author, who facilitated every user study session, created an initial codebook by reviewing the revised study transcripts. The research team then collaboratively iterated on the codebook while checking for bias and coverage. With a final codebook consisting of 23 codes, the first author coded participants’ quotes, after which the team discussed the resulting themes. For likert-scale score analysis, we used a Wilcoxon signed-rank test since the data does not follow a normal distribution.
@@ 6. Results
Effectiveness. In examining Likert data on perceived effectiveness, the findings showed no significant difference (; ). However, participants on average gave higher ratings to affordance augmentations (5.3; 1.6) over whole object augmentations (4.6; 1.4). Affordance augmentations are advantageous in quickly understanding the overall scene (P1, P10), along with the placement and orientation of individual objects: “It helps to have two colors. I could see that more readily and quickly to understand how to use the object and how it is placed… your system is helpful because where the tool starts and ends and where the handle starts and ends is more clear” (P5). In addition, affordance augmentations become particularly useful when handling objects that have hazardous (9 out of 10 participants) (e.g., sharp, hot ) or small (8/10) (e.g., door handles, buttons on appliances) parts, or have insufficient color contrast (7/10) (e.g., all silver or black cooking tools): “You’ve gotta show parts you can and shouldn’t grab. Green tells me that’s a safe place to go with my hand. Anything not green, I shouldn’t grab… It can help me avoid dangerous parts or perhaps even find small things like remote controllers” (P4). Furthermore, four participants expressed that for some objects with more complex interaction components than “just grab and don’t grab” (P7), like a carafe with its handle, base, buttons, lid, and spout, they would accept the use of more than two colors, although “more than four colors can be quite distracting” (P7). We discuss augmentation designs in Section 6.3.
@@
Distraction & Comfort. While participants qualitatively expressed that the whole object augmentations are less distracting than the affordance augmentations (6/10), the comfort and distraction Likert data was not significantly different (; and ; respectively). The difference in average rating was also negligeable, although participants on average found whole object augmentations to be slightly more comfortable (5.1; 1.3 vs. 5.0; 1.2) and less distracting (2.3; 0.9 vs. 2.5; 1.1). P6, who preferred whole object augmentations during the study, said “I think the more colors you have, the more distracting it becomes. So I prefer just the whole object in green than having 2 or 3 different colors. An outline would be better. I would definitely stay away from multicolored and just stick with one color. I can figure out its different parts.” Additionally, three participants shared that the whole object augmentations could be more useful depending on the scenarios. For tasks such as locating or avoiding objects, where interaction is not the goal, whole object augmentations are more preferable, since they are less distracting: “If I am looking for the remote controller, if it could make the remote stand more out in green or something. I don’t need its parts” (P3).
In summary, if a person’s intent is to interact with an object, affordance augmentations are more helpful than whole object augmentations. Conversely, in cases where interaction is not the objective, whole object augmentations may be preferred as they are less distracting. In our study, participants had to interact with various kitchen tools, hence most of them found our affordance augmentations more useful.
6.2. Experiences with CookAR
@@
All participants were able to complete all free-form cooking steps within three to five minutes. However, due to current technical limitations in accurately segmenting object affordances, deploying heavy CV models, and rendering spatially accurate overlays in real-time on AR headsets, CookAR experienced recognition errors and perceived latency, which affected LV users’ experiences. All participants observed “flickering” and inaccurate affordance augmentations. Participants also pointed out that “the colors took some time to catch up” (P3) as they quickly rotated their head. Nonetheless, all but P6 saw potential in CookAR to assist with kitchen tool interactions and beyond: “I like the contrasting color. I just wish it more closely matched the object’s actual location. I think this highlighting scene is a great start. If the system is perfect, the dual color highlight system would be great and most useful. The system would be perfect if I am in a kitchen or just trying to grab really anything” (P1). P1, P5, P9, and P10 were particularly excited as they were able to better visually perceive object information: “This is fun! I can also use my eyes more to see shape and how it can be used. I want to try your system again in the future once you make it better” (P9).
@@
Recognizing the potential of CookAR, participants identified additional use cases of a CookAR-like system that go beyond a cooking scenario, including cleaning (P3, P6, P7, P9), woodworking (P4, P5, P9), walking outdoors (P2, P6, P9), driving (P4, P6, P7), visiting a foreign kitchen (P2, P5), restaurants (P3, P9), gardening (P5, P10), watching sports (P7, P10), playing board games (P7, P10), going down stairs (P7, P9), identifying pill bottles (P7), and interacting with appliances with multiple buttons like a toaster (P1).
@@ 6.3. Desired Augmentation Designs
Combining solid and outline augmentations. As opposed to solid-colored overlays, nine participants preferred a mix of solid and outline augmentations, because solid colors are more salient, whereas outlines are less distracting: “Solid colors are helpful because they grab my attention… outlines are helpful because I can still see the part I’m trying to use with less distraction” (P5). Among those nine participants, all but P7 preferred solid-colored overlays for the grabbable area and outlines for the hazardous area because “the grabbable area is the most important” (P3, P4, P5, P8, P9), “all you need to know is its shape” (P4, P9), and “other parts should be outlined since you may want to do more with it, and solid color just makes it harder to use it” (P4, P8, P9). However, P7 preferred the outline for the grabbable area since it is not distracting and still shows the shape of the handle.
For the risky area to outline, P8 preferred highlighting solely the exact hazard (e.g., the sharp edge of the knife blade), as opposed to the entire dangerous part of an object (e.g., the whole knife blade), because she needs to know the relatively safer area for interaction. For example, as she described, “I might grab the top of the blade when I’m dicing or chopping. This tells you exactly where you shouldn’t touch.” In contrast, all other participants wanted the outline augmentation because it is less distracting, yet still defines the overall shape of the hazardous area (P4, P5, P7, P9, P10) and what it is used for (P4, P7, P9). For example, P9 said, “I prefer to see the outlines on the [whole] blade, just so that way, you know which type of blade you’re grabbing. Cause a bread knife would look different from your knife. Some are thinner, some are fatter. People can be quite picky about their knife choices.” P3, who preferred to outline dangerous parts, cautioned that “depending on how low vision you are on the spectrum, you may need solid colors to be able to see some of these objects.”
@@
Every participant favored using green for safe-to-grab and red for dangerous areas, as “green signals ‘yes’ while ‘red’ signals no” (P4). However, P4, P7, and P9 struggled to clearly see our choice of red and requested a brighter shade of red, with P4 even suggesting white. Moreover, participants noted that the color contrast between the tool and the background is more important than the specific colors used, since a lot of objects in a kitchen are white, silver, or black with low contrast. For example, when cooking mac and cheese in the study, most (8/10) participants found it challenging to cut butter and understand where the yellow butter starts and ends because it was on a yellow wooden cutting board. To address this, P3, P4, P5, and P10 suggested the system should automatically select colors that contrast against the background: “The background you have it against will make a big difference, right? So on a darker background, I should be getting light colors” (P4). P7 jokingly said, “I mean, a green stick of butter could be weird, but it would at least let me cut better.”
Auditory feedback. Instead of visual augmentations, all participants preferred auditory feedback for warning in urgent scenarios (e.g., when the user’s hands get too close to a knife blade), as a visual warning could be easily missed by low vision users (P3, P5, P7, P9) and also makes the overall visual field busier (P4, P10). P3, P5, and P6 suggested small yet noticeable audio such as “beep beep,” while P7 and P10 preferred explicit verbal waning (e.g., “stop”) since small noises can also be generated by other devices, such as a microwave or a fridge. P4 and P9 further suggested the system should employ different auditory signals for different hazards.
@@ 6.4. Additional Tool Affordances
In addition to the grabbable and hazardous affordances focused in our CookAR system, participants collectively suggested five other important affordances for kitchen tools: (1) entry area, (2) exit area, (3) containment area, (4) intersection area, and (5) activation area. We elaborate on these seven affordances along with participants’ preferred augmentation designs.
Grabbable area. A grabbable area is the part designed for safe handling or manipulation. This can include handles, grips, or any part intended for direct hand contact. For grabbable part of an object, participants preferred green solid-colored augmentations (Section 6.3).
Hazardous area. A hazard area is the part that poses potential risks or dangers to the user. This could include sharp edges, hot surfaces, or any part that can cause injury if touched or mishandled. For hazardous part of an object, participants preferred red outline augmentations (Section 6.3).
Entry area. An entry area is the part designed for initiating access, such as pouring. This could be the rim of a cup or a pot, the opening of a carafe, or any designated point that allows entry into an object’s containment space. All participants consistently noted that this area on an object should be augmented by an outline rather than a solid color, as the latter obstructs relevant actions like pouring or scooping: “The color blobs hide the item that you’re trying to put things into virtually completely. And so I can’t really tell if I am pouring something in correctly” (P4).
@@
Additionally, eight participants expressed that they need assistance with understanding the depth of the containment space and the amount of content it already holds. As P8 expressed, “A lot of people with low vision cannot see inside and know how much water they can pour. So somehow showing the water level and size of the teapot [is helpful]. Mine is a lot bigger, it makes 12 cups or something, and it’s all black, so it’s even harder to see what’s inside.” While P9 has a strategy to overcome this challenge by using her finger to feel the liquid level, she cannot use it when the water is hot. She thus suggested the system rendering “a blue disk” to indicate the water level. In terms of augmenting the depth of the containment area, participants suggested using a virtual line from the rim to the bottom of the pot (P3, P4, P5, P9, P10), a measuring tape with ticks (P4, P5, P9, P10), or a line with changing colors (e.g., a green line that turns red as water fills up), (P4, P5). P10 further suggested an auditory cue (e.g., a ‘ding’ sound) to indicate action milestones, such as when reached quarter of a cup.
@@
Activation area. An activation area is designed for initiating, activating, or turning on an object’s function or features. This could be buttons, switches, touch-sensitive surfaces, or any interactive components that trigger the operation of an object. Participants identified activation areas on many household appliances, such as buttons and dials on stove tops, microwaves, or coffee pots. They are used for various purposes including starting a machine, opening a lid, and adjusting settings. Seven participants preferred outline augmentations for the activation area: “I just bought a vacuum with multiple buttons. You would want different colors for the handles and buttons” (P8). P10 further suggested a clock-like augmentation in addition to the outline for turnable dials: “On a stove, I don’t know what is medium heat. As I turn the knobs on a stove, the system could show me ‘2 o’clock,’ ‘3 o’clock,’ and so on. ‘6 o’ clock’ is probably a medium heat. ‘9 o’clock’ is probably a high heat.”
@@ 7. Discussion
CookAR explores the use of affordance augmentations to enhance kitchen tool interactions for LV people and advances the state-of-the-art in AI-powered AR systems. Results from our user study indicates a preference for affordance augmentations over whole object during tool interactions. Additionally, participants favored augmentation designs that incorporate both solid-colored and outlined overlays with contrasting colors. We discuss design implications for affordance augmentations as well as current limitations and future opportunities of AI-powered AR systems for low vision.
@@ 7.1. Design Implications for Affordances
Throughout the study, our LV participants proposed a range of affordances for kitchen tools and indicated their preferred augmentation designs for each. In this section, we summarize and expand on these suggestions.
@@
When to use affordance augmentations? Our study indicates that visual augmentations should maximize utility and minimize visual clutter and confusion. As such, it is critical to render augmentations tailored to users’ intent and reduce distraction. For example, affordance augmentations that involve multiple pieces and colors would be more preferred to support interactions, while basic whole-object augmentations are more suitable in general visual perception tasks such as avoiding obstacles. Beyond a kitchen, affordance augmentations could also be applied to other scenarios, as our participants suggested, such as gardening, playing board games, and interacting with appliances (reaffirming (Lang and Machulla, 2021)).
@@
Where to apply affordance augmentations? Affordances can refer to any object parts that indicate diverse action or interaction opportunities. However, LV people face distinct interaction challenges, resulting in unique affordances to augment for them. In our qualitative study, we identified seven essential affordances of kitchen tools that represent important yet challenging interaction tasks for LV users. They include: (1) grabbable area, affording touching and handling action; (2) hazardous area, affording risks and avoidance; (3) entry, affording a target to aim at or pour in; (4) exit, affording pouring out and usually requiring accurate alignment with the entry of another object (e.g., food transferring or pouring); (5) containment area, affording holding content in and preferring augmentations on the content amount (e.g., ingredient measurement); (6) intersection area, affording touching or interaction between two objects; and (7) activation area, affording control features on an object. This affordance framework summarizes the critical areas on objects as well as the hand-object (e.g., grabbable vs. hazardous areas) and object-object (e.g., entry-exit alignment, intersection between objects) relationships during interactions.
@@
How to augment affordances? Different augmentations should be designed for different types of affordances according to the interaction tasks they indicate. Our study revealed the most preferred augmentations for different affordances. Specifically, solid-colored overlays with great visual emphasis are preferred for grabbable area to enable fast perception and action, measuring augmentations (e.g., line with ticks) for containment area to indicate content amount, and outlines for other affordances to avoid distraction and occlusion. In terms of colors, augmentations should adopt colors with high contrast against the environment. We also suggest leveraging the commonly-agreed meanings of colors, such as green for safe-to-grab areas and red for risky areas. However, while preferred by the LV participants in our study, the green-red combination should be used cautiously considering that LV users may experience color deficiency.
@@
How to control affordance augmentations? Due to the diverse visual abilities and preferences of LV users (Zhao et al., 2015), future systems should support extensive personalization capabilities such as voice-based control for customization and automatic adaptations. For instance, users should be able to adjust different aspects of an augmentation’s design, such as switching it on or off, choosing between solid-colored and outlined overlays, changing the outline’s thickness, and selecting suitable colors. Additionally, these systems could intelligently adapt by recognizing user actions or tasks to only highlight necessary object parts and signal warnings. They can also automatically alter the augmentations’ colors to generate high contrast against the background.

### v2

@@ 5.4. Analysis
We recorded participants’ quotes using Zoom. Transcriptions were first done by the video conferencing software, then the research team manually revised the transcripts. We collected 346 distinct quotes across our 10 LV participants, which we analyzed using reflexive thematic analysis (Braun and Clarke, 2006, 2019). The first author, who facilitated every user study session, created an initial codebook by reviewing the revised study transcripts. The research team then collaboratively iterated on the codebook while checking for bias and coverage. With a final codebook consisting of 23 codes, the first author coded participants’ quotes, after which the team discussed the resulting themes. For Likert score analysis, we used a Wilcoxon signed-rank test since the data does not follow a normal distribution.
@@ 6. Results
Effectiveness. In examining Likert scores on perceived effectiveness, we found no significant difference between whole object augmentations and affordance augmentations (; ). However, participants on average gave higher ratings to affordance augmentations (; ) over whole object augmentations (; ). Affordance augmentations are advantageous in quickly understanding the overall scene (P1, P10), along with the placement and orientation of individual objects: “It helps to have two colors. I could see that more readily and quickly to understand how to use the object and how it is placed… your system is helpful because where the tool starts and ends and where the handle starts and ends is more clear” (P5). In addition, affordance augmentations become particularly useful when handling objects that have hazardous (9 out of 10 participants) (e.g., sharp, hot ) or small (8/10) (e.g., door handles, buttons on appliances) parts, or have insufficient color contrast (7/10) (e.g., all silver or black cooking tools): “You’ve gotta show parts you can and shouldn’t grab. Green tells me that’s a safe place to go with my hand. Anything not green, I shouldn’t grab… It can help me avoid dangerous parts or perhaps even find small things like remote controllers” (P4). Furthermore, four participants expressed that for objects with more complex interaction components than “just grab and don’t grab” (P7), like a carafe with its handle, base, buttons, lid, and spout, they would accept the use of more than two colors, although “more than four colors can be quite distracting” (P7). We discuss additional augmentation designs in Section 6.3.
@@
Distraction & Comfort. While most participants qualitatively expressed that the whole object augmentations are less distracting than the affordance augmentations (6/10), the comfort and distraction Likert scores were not significantly different (; and ; respectively). The difference in average rating was also negligible, although participants on average found whole object augmentations to be slightly more comfortable (; vs. ; ) and less distracting (; vs. ; ). P6, who preferred whole object augmentations, said “I think the more colors you have, the more distracting it becomes. So I prefer just the whole object in green than having 2 or 3 different colors. An outline would be better. I would definitely stay away from multicolored and just stick with one color. I can figure out its different parts.” Additionally, three participants shared that the whole object augmentations could be more useful depending on the scenarios. For tasks such as locating or avoiding objects, where interaction is not the goal, whole object augmentations are more preferable, since they are less distracting: “If I am looking for the remote controller, if it could make the remote stand more out in green or something. I don’t need its parts” (P3).
In summary, we found that if a person intends to interact with an object, affordance augmentations are more helpful than whole object augmentations. Conversely, in cases where interaction is not the objective, whole object augmentations may be preferred as they are less distracting. Despite the positive qualitative feedback on CookAR, the Likert scores for the real-world baseline (without CookAR) were higher than those for both the whole object augmentations (effectiveness: ; , comfort: ; , ; distraction: ) and affordance augmentations (effectiveness: ; , comfort: ; , distraction: ; ). This may be due to current limitations of AI-powered AR systems such as accuracy and latency. We discuss this further in Section 7.2.
6.2. Free-form Cooking with CookAR
@@
All participants were able to complete all free-form cooking steps within three to five minutes. However, due to technical limitations in accurately segmenting affordances, deploying heavy CV models, and rendering spatially accurate overlays in real-time on AR headsets, participants experienced recognition errors and latency with CookAR. All participants observed “flickering” and inaccurate augmentations. Participants also pointed out that “the colors took some time to catch up” (P3) as they quickly rotated their head.
@@
Nonetheless, all but P6 saw potential in CookAR to assist with kitchen tool interactions and beyond: “I like the contrasting color. I just wish it more closely matched the object’s actual location. I think this highlighting scene is a great start. If the system is perfect, the dual color highlight system would be great and most useful. The system would be perfect if I am in a kitchen or just trying to grab really anything” (P1). P1, P5, P9, and P10 were particularly excited as they were able to better visually perceive object information: “This is fun! I can also use my eyes more to see shape and how [a tool] can be used. I want to try your system again once you make it better” (P9). Participants also identified the following additional use cases for a CookAR-like system: cleaning (P3, P6, P7, P9), woodworking (P4, P5, P9), walking outdoors (P2, P6, P9), driving (P4, P6, P7), visiting a foreign kitchen (P2, P5), restaurants (P3, P9), gardening (P5, P10), watching sports (P7, P10), playing board games (P7, P10), going down stairs (P7, P9), identifying pill bottles (P7), and interacting with appliances with multiple buttons like a toaster (P1).
@@ 6.3. Desired Augmentation Designs
Combining solid and outline augmentations. As opposed to solid-colored overlays, nine participants preferred a mix of solid and outline augmentations because solid colors are more salient, whereas outlines are less distracting: “Solid colors are helpful because they grab my attention… outlines are helpful because I can still see the part I’m trying to use with less distraction” (P5). Among those nine participants, all but P7 preferred solid-colored overlays for the grabbable area and outlines for the hazardous area because “the grabbable area is the most important” (P3, P4, P5, P8, P9), “all you need to know is its shape” (P4, P9), and “other parts should be outlined since you may want to do more with it, and solid color just makes it harder to use it” (P4, P8, P9). However, P7 preferred the outline for the grabbable area since it is less distracting and still shows the shape of the handle.
For the risky area to outline, P8 preferred highlighting solely the exact hazard (e.g., the sharp edge of the knife blade), as opposed to the entire dangerous part of an object (e.g., the whole knife blade), because she needs to know the relatively safer area for interaction. For example, as she described, “I might grab the top of the blade when I’m dicing or chopping. This tells you exactly where you shouldn’t touch.” In contrast, all other participants wanted the outline augmentation because it is less distracting, yet still defines the overall shape of the hazardous area (P4, P5, P7, P9, P10) and what it is used for (P4, P7, P9). For example, P9 said, “I prefer to see the outlines on the [whole] blade, just so that way, you know which type of blade you’re grabbing. Cause a bread knife would look different from your knife. Some are thinner, some are fatter. People can be quite picky about their knife choices.”
@@
Every participant favored using green for safe-to-grab and red for dangerous areas, as “green signals ‘yes’ while ‘red’ signals no” (P4). However, P4, P7, and P9 struggled to clearly see our choice of red and requested a brighter shade of red, with P4 even suggesting white. Moreover, participants noted that the color contrast between the tool and the background is more important than the specific colors used, since many kitchen objects are white, silver, or black with low contrast. For example, when cooking mac and cheese in the study, most (8/10) participants found it challenging to cut butter and understand where the yellow butter starts and ends because it was on a yellow wooden cutting board. To address this, P3, P4, P5, and P10 suggested the system should automatically select colors that contrast against the background: “The background you have it against will make a big difference, right? So on a darker background, I should be getting light colors” (P4). P7 jokingly said, “I mean, a green stick of butter could be weird, but it would at least let me cut better.”
Auditory feedback. Instead of visual augmentations, all participants preferred auditory feedback for warning in urgent scenarios (e.g., when the user’s hands get too close to a knife blade), as a visual warning could be easily missed by low vision users (P3, P5, P7, P9) and also makes the overall visual field busier (P4, P10). P3, P5, and P6 suggested short yet noticeable audio such as “beep beep,” while P7 and P10 preferred explicit verbal warning (e.g., “stop”) since small noises can also be generated by other devices, such as a microwave or a fridge. P4 and P9 further suggested the system should employ different auditory signals for different hazards.
@@ 6.4. Additional Tool Affordances
In addition to the grabbable and hazardous affordances focused on by our CookAR system, participants collectively suggested five other important affordances for kitchen tools: (1) entry area, (2) exit area, (3) containment area, (4) intersection area, and (5) activation area. We elaborate on these seven affordances along with participants’ preferred augmentation designs.
Grabbable area. A grabbable area is the part designed for safe handling or manipulation. This can include handles, grips, or any part intended for direct hand contact. For grabbable part of an object, participants preferred green solid-colored augmentations.
Hazardous area. A hazard area is the part that poses potential risks or dangers to the user. This could include sharp edges, hot surfaces, or any part that can cause injury if touched or mishandled. For hazardous part of an object, participants preferred red outline augmentations.
Entry area. An entry area is the part designed for initiating access, such as pouring. This could be the rim of a cup or a pot, the opening of a carafe, or any designated point that allows entry into an object’s containment space. All participants consistently noted that this area should be augmented by an outline rather than a solid color, as the latter obstructs relevant actions like pouring or scooping: “The color blobs hide the item that you’re trying to put things into virtually completely. And so I can’t really tell if I am pouring something in correctly” (P4).
@@
Additionally, eight participants expressed that they need assistance with understanding the depth of the containment space and the amount of content it already holds. As P8 expressed, “A lot of people with low vision cannot see inside and know how much water they can pour. So somehow showing the water level and size of the teapot [is helpful]. Mine is a lot bigger, it makes 12 cups or something, and it’s all black, so it’s even harder to see what’s inside.” While P9 has a strategy to overcome this challenge by using her finger to feel the liquid level, she cannot use it when the water is hot. She thus suggested the system rendering “a blue disk” to indicate the water level. In terms of augmenting the depth of the containment area, participants suggested using a virtual line from the rim to the bottom of the pot (P3, P4, P5, P9, P10), a measuring tape with ticks (P4, P5, P9, P10), or a line with changing colors (e.g., a green line that turns red as water fills up), (P4, P5). P10 further suggested an auditory cue (e.g., a ‘ding’ sound) to indicate action milestones, such as when water reaches quarter of a cup.
@@
Activation area. An activation area is designed for initiating, activating, or turning on an object’s function or features. This could be buttons, switches, touch-sensitive surfaces, or any interactive components that trigger the operation of an object. Participants identified activation areas on many household appliances, such as buttons and dials on stove tops, microwaves, or coffee pots. They are used for various purposes including starting a machine, opening a lid, and adjusting settings. For example, P8 said: “I just bought a vacuum with multiple buttons. You would want different colors for the handles and buttons” Seven participants preferred outline augmentations for the activation area. Additionally, P10 further suggested a clock-like augmentation in addition to an outline for turnable dials: “On a stove, I don’t know what is medium heat. As I turn the knobs on a stove, the system could show me ‘2 o’clock,’ ‘3 o’clock,’ and so on. ‘6 o’ clock’ is probably a medium heat. ‘9 o’clock’ is probably a high heat.”
@@ 7. Discussion
CookAR explores the use of real-time affordance augmentations to enhance kitchen tool interactions for LV people and advances the state-of-the-art in AI-powered AR systems. Results from our user study indicates a preference for affordance augmentations over whole object augmentations during tool interactions. Additionally, participants favored augmentation designs that incorporate both solid-colored and outlined overlays with contrasting colors. We discuss design implications for affordance augmentations as well as current limitations and future opportunities of AI-powered AR systems for low vision.
@@ 7.1. Design Implications for Affordances
Throughout the study, LV participants proposed a range of affordances for kitchen tools and indicated preferred augmentation designs. We summarize and expand on these suggestions.
@@
When to use affordance augmentations? Our study findings suggest that visual augmentations should maximize utility and minimize visual clutter and confusion. As such, it is critical to render augmentations tailored to users’ intent and reduce distraction. For example, affordance augmentations that involve multiple pieces and colors are more preferred to support direct hand-object interactions, while whole object augmentations are more suitable in general visual perception tasks such as avoiding obstacles and locating an object. Beyond a kitchen, affordance augmentations could also be applied to other scenarios, as our participants suggested, such as gardening, playing board games, and interacting with appliances (reaffirming Lang et al. (Lang and Machulla, 2021)).
@@
Where to apply affordance augmentations? Affordances can refer to any object parts that indicate diverse actions or interaction opportunities. However, LV people face distinct interaction challenges, resulting in unique affordance opportunities. In our qualitative study, we identified seven essential affordances of kitchen tools that encapsulate important yet challenging interaction tasks for LV users. They include: (1) grabbable area, affording touching and handling action; (2) hazardous area, affording risks and avoidance; (3) entry, affording a target to aim at or pour in; (4) exit, affording pouring out and usually requiring accurate alignment with the entry of another object (e.g., food transferring or pouring); (5) containment area, affording holding content in and preferring augmentations on the content amount (e.g., ingredient measurement); (6) intersection area, affording touching or interaction between two objects; and (7) activation area, affording control features on an object. This affordance taxonomy summarizes the critical areas on objects as well as the hand-object (e.g., grabbable vs. hazardous areas) and object-object (e.g., entry-exit alignment, intersection between objects) relationships during interactions.
@@
How to augment affordances? Different augmentations should be designed for different affordances according to the interaction tasks they indicate. In our study, participants preferred solid-colored overlays for grabbable areas to enable fast perception and action, measuring augmentations (e.g., line with ticks) for containment area to indicate content amount, and outlines for other affordances to avoid distraction and occlusion. In terms of colors, augmentations should adopt colors with high contrast against the environment. We also suggest leveraging cultural and semantic meanings of colors, such as green for safe-to-grab areas and red for risky areas. However, while preferred by the LV participants in our study, the green-red combination should be used cautiously given the prominence of red-green color vision deficiency.
@@
How to control affordance augmentations? Due to the diverse visual abilities and preferences of LV users (Zhao et al., 2015), future systems should support extensive personalization capabilities such as voice-based control for customization and automatic adaptations. For instance, users should be able to adjust different aspects of an augmentation, such as switching it on or off, choosing between solid-colored and outlined overlays, changing the outline thickness, and selecting suitable colors. Additionally, these systems should intelligently adapt by recognizing user actions or tasks to only highlight necessary object parts and signal warnings. They should also automatically alter the augmentation colors to generate high contrast against the background.
Example designs. Reflecting on these key design insights for affordance augmentations, we created some initial design mockups shown in Figure 7. Starting with solid-colored overlays for grabbable areas, measuring lines for containment areas, and outlines for all other areas, CookAR should allow customization, such as toggling overlays, adjusting outline thickness, and changing colors. Future work should explore these and other designs empirically.